Page headers and footers in L^{A}TEX

Piet van Oostrum^{∗} Dept. of Computer Science Utrecht University

March 5, 1996

Abstract

Standard L^{A}TEX offers the page styles empty, plain, headings and myheadings. For simple page layouts these suffice, but they are rather rigid. A more flexible page layout may be obtained by using the fancyheadings package. This article describes how to create advanced page headers and footers in L^{A}TEX documents by using this package. Some of the techniques described are also applicable with other page styles.

Contents

1 What is fancyheadings . . . . . . . . . . . . . . . . . . . . . . . . 2 2 Where to get it? . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 Use of fancyheadings . . . . . . . . . . . . . . . . . . . . . . . . . 2 4 A simple example . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 5 An example of two-sided printing . . . . . . . . . . . . . . . . . . 4 6 Redefining plain style . . . . . . . . . . . . . . . . . . . . . . . . 4 7 The default layout . . . . . . . . . . . . . . . . . . . . . . . . . . 6 8 The scoop on L^{A}TEX’s marks . . . . . . . . . . . . . . . . . . . . . 7 9 Dictionary style headers . . . . . . . . . . . . . . . . . . . . . . . 9 10 Fancy layouts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 11 Two book examples . . . . . . . . . . . . . . . . . . . . . . . . . 11 12 Special page layout for float pages . . . . . . . . . . . . . . . . . 13 13 Those blank pages . . . . . . . . . . . . . . . . . . . . . . . . . . 13 14 N of M style page numbers . . . . . . . . . . . . . . . . . . . . . 14 15 When to change the headers and footers? . . . . . . . . . . . . . 15 16 Headers and footers induced by the text . . . . . . . . . . . . . . 16 17 A movie . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 18 Thumb-indexes . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 19 Final Remarks . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

∗A considerable part of this article was written by George Grätzer (University of Manitoba)

in Notices Amer. Math. Soc. Thanks, George!

1 What is fancyheadings

The fancyheadings macro package allows you to customize in L^{A}TEX your page headers and footers in an easy way. You can define:

• three-part headers and footers

• decorative lines in headers and footers

• headers and footers wider than the width of the text

• multi-line headers and footers

• separate headers and footers for even and odd pages

• different headers and footers for chapter pages

Of course, you also have complete control over fonts, uppercase and lowercase displays, etc.

2 Where to get it?

You only need the file fancyheadings.sty which you can find at the CTAN

sites: ftp.shsu.edu (U.S.), ftp.tex.ac.uk (U.K.), and ftp.dante.de (Ger-

many). You will find it in the directory

/tex-archive/macros/latex209/contrib/fancyheadings. Although fancy-

headings is a L^{A}TEX 2.09 style file, it will work with L^{A}TEX 2ε. Also this article and the two packages fixmarks and extramarks can be found there. (The code for these packages given in this article is a simplified version.)

3 Use of fancyheadings

To use this package in a L^{A}TEX 2ε document, place the file fancyheadings.sty in a directory/folder where TEX can find it (normally in the input directory/folder), and include in the preamble of your document after

\documentclass{...}

the commands:

\usepackage{fancyheadings}
\pagestyle{fancy}

We can visualize the page layout we can create with fancyheadings as follows:

LeftHeader  CenteredHeader  RightHeader

page body

LeftFooter  CenteredFooter  RightFooter

The LeftHeader and LeftFooter are left justified; the CenteredHeader and CenteredFooter are centered; the RightHeader and RightFooter are right justified.

We define each of the six “fields” and the two decorative lines separately.

4 A simple example

K. Grant is writing a report to Dean A. Smith, on “The performance of new graduates” with the following page layout:

The performance of new graduates

page body

From: K. Grant  To: Dean A. Smith  3

where “3” is the page number. The title: “The performance of new graduates” is bold.

This is accomplished by these commands following \pagestyle{fancy}:

\lhead{}
\chead{}
\rhead{\bfseries The performance of new graduates}
\lfoot{From: K. Grant}
\cfoot{To: Dean A. Smith}
\rfoot{\thepage}
\setlength{\headrulewidth}{0.4pt}
\setlength{\footrulewidth}{0.4pt}

(The \thepage macro displays the current page number. \bfseries is the L^{A}TEX 2ε’s way of selecting bold face.)

This is now fine, except that the first page does not need all these headers and footers. To eliminate all but the centered page number, issue the command

\thispagestyle{plain}
after the \begin{document} and the \maketitle commands.

Alternatively, issue

\thispagestyle{empty}

if you do not want any headers or footers.

In fact the standard L^{A}TEX classes have the command \maketitle defined in such a way that a \thispagestyle{plain} is automatically issued. So if you do want the fancy layout on a page containing \maketitle you must issue

a \thispagestyle{fancy} after the \maketitle.

5 An example of two-sided printing

Some document classes, such as book.cls, print two-sided: the even pages and the odd pages have different layouts; other document classes use the twoside option to print two-sided.

Now let us print the report two-sided. Let the above page layout be used for the odd (right-side) pages, and the following for the even (left-side) pages:

The performance of new graduates

page body

4  From: K. Grant  To: Dean A. Smith

where “4” is the page number.

Here are the commands:

\lhead[]{\bfseries The performance of new graduates}
\chead{}
\rhead[\bfseries The performance of new graduates]{}
\lfoot[\thepage]{From: K. Grant}
\cfoot[From: K. Grant]{To: Dean A. Smith}
\rfoot[To: Dean A. Smith]{\thepage}
\setlength{\headrulewidth}{0.4pt}
\setlength{\footrulewidth}{0.4pt}

As you see, the values of the fields for the even (left-side) pages are given in the square brackets. Again, you may use \thispagestyle{plain} for a simple page layout for page 1.

6 Redefining plain style

Some L^{A}TEX commands, like \chapter, use the \thispagestyle command to automatically switch to the plain page style, thus ignoring the page style currently in effect. To customize even such pages use the fancyplain page style. This page style sets up fancy for normal pages and in addition redefines the page style plain to conform more to the page style fancy. For instance, if all the page numbers are bold, you may want this on the first page also. You do this with the command

\pagestyle{fancyplain}
Note that this does not issue a \pagestyle{plain} or \thispagestyle{plain}

command. This is either done by L^{A}TEX (for instance, for the first page of a chapter in the book class), or you have to do it explicitly with the

\thispagestyle{plain}

command.

The redefinition of the plain style is done with the command

\fancyplain{plain_value}{normal_value}

inserted into all the appropriate arguments. As an example, let us redefine the plain style for the report in Section 5 by making the page number bold.

The leftheader definition:

\lhead[]{\bfseries The performance of new graduates}

changes to

\lhead[\fancyplain{}{}]{\fancyplain{}{\bfseries The
   performance of new graduates}}

Note that the first argument of both \fancyplain commands is empty, since the redefined plain style has an empty leftheader. (The \fancyplain{}{} is superfluous but for symmetry reasons it is given in this example.)

The left- and rightfooter definitions change more. For instance, the leftfooter definition changes from

\lfoot[\thepage]{From: K. Grant}

to

\lfoot[\fancyplain{}{\bfseries \thepage}]
      {\fancyplain{}{From: K. Grant}}

Carrying out these changes for all six fields, we get the code:

\lhead[\fancyplain{}{}]{\fancyplain{}{\bfseries The
   performance of new graduates}}
\chead{\fancyplain{}{}}
\rhead[\fancyplain{}{\bfseries The performance of new
   graduates}]{\fancyplain{}{}}
\lfoot[\fancyplain{}{\thepage}]{\fancyplain{}{From: K. Grant}}
\cfoot[\fancyplain{\bfseries \thepage}{From: K.
   Grant}]{\fancyplain{\bfseries \thepage}{To: Dean A. Smith}}
\rfoot[\fancyplain{}{To: Dean A. Smith}]{\fancyplain{}%
   {\bfseries \thepage}}
\setlength{\headrulewidth}{0.4pt}
\setlength{\footrulewidth}{0.4pt}

For the “fancy plain” page style, the thickness of the lines is defined by

            \plainheadrulewidth and \plainfootrulewidth

(both default to 0pt). Observe that you may have different “fancy plain” page styles for even and odd pages.

To use the “fancy plain” page style, define these macros; in the preamble, include

\usepackage{fancyheadings}
\pagestyle{fancyplain}

and invoke the page style “fancy plain” with

\thispagestyle{plain}

where desired.

7 The default layout

Let us use the book.cls documentclass and the default settings for fancyheadings; so we only issue the commands

\usepackage{fancyheadings}
\pagestyle{fancy}

and let fancyheadings take care of everything. On the pages where new chapters start, we get a centered page number in the footer; there is no header, and there are no decorative lines.

On an even page, we get the layout:

1.2 EVALUATION CHAPTER 1. INTRODUCTION

page body

4

On an odd page, we get the layout:

CHAPTER 1. INTRODUCTION 1.2 EVALUATION

page body

3

where the header text is slanted uppercase.

This default layout is produced by the following commands:

\lhead[\fancyplain{}{\slshape \rightmark}]{\fancyplain{}%
   {\slshape \leftmark}}
\chead{}
\rhead[\fancyplain{}{\slshape \leftmark}]{\fancyplain{}%
   {\slshape \rightmark}}
\lfoot{}
\cfoot{\rmfamily \thepage}
\rfoot{}

The following settings are used for the decorative lines:

 \headrulewidth  0.4pt
 \footrulewidth  0 pt
 \plainheadrulewidth  0 pt
 \plainfootrulewidth  0 pt

The header text is turned into all uppercase in book.cls.

8 The scoop on L^{A}TEX’s marks

Usually, for documents of class book and report, you may want to use chapter and section information in the headings (chapter only for one-sided printing), and for documents of class article, section and subsection information (section only for one-sided printing). L^{A}TEX uses a marker mechanism to remember the chapter and section (section and subsection) information for a page; this is discussed in detail in the L^{A}TEX Companion, Section 4.3.1.

There are two ways you can use and change the higher- and lower-level sectioning information available to you. The macros: \leftmark (higher-level) and \rightmark (lower-level) contain the information processed by L^{A}TEX, and you can use them directly as shown in Section 7.

The \leftmark contains the left argument of the last \markboth on the page, the \rightmark contains the right argument of the first \markboth or the only argument of the first \markright on the page. If no marks are present on a page they are “inherited” from the previous page.

You can influence how chapter, section, and subsection information (only two of them!) is displayed by redefining the \chaptermark, \sectionmark, and \subsectionmark commands^{1}. You must put the redefinition after the first call of \pagestyle{fancy} as this sets up the defaults.

Let us illustrate this with chapter info. It is made up of three parts:

• the number (say, 2), displayed by the macro \thechapter

• the name (in English, Chapter), displayed by the macro \chaptername

• the title, contained in the first argument of \markboth.

Figure 1 shows some variants for “Chapter 2. Do it now” (the last example is appropriate in some non-English languages).

For the lower-level sectioning information, do the same with \markright. So if “Section 2.2. First steps” is the current section, then

\renewcommand{\sectionmark}[1]{
   \markright{\thesection.\ #1}}

will give “2.2. First steps”

Redefining the \chaptermark and \sectionmark commands may not eliminate all uppercaseness. E.g. the bibliography will have a title of bibliography in the header, as the \uppercase is explicitly given in the definition of \thebibliography. Similar for index etc. If you don’t want to redefine these commands, you can use an ugly hack like the following:

\lhead{\let\uppercase\relax\rightmark}
\rhead{\let\uppercase\relax\leftmark}

It should be noted that the L^{A}TEX marking mechanism works fine with chapters (which always start on a new page) and sections (which are reasonably long). It does not work quite as well with short sections and subsections. This is a problem with L^{A}TEX, not with fancyheadings.

1There are similar commands for paragraph and subparagraph but they are seldom used.

Figure 1 Marker variants

Code:

\renewcommand{\chaptermark}[1]{
 \markboth{\chaptername\
 \thechapter.\ #1}{}}
\renewcommand{\chaptermark}[1]{
 \markboth{\uppercase{
 \chaptername}\ \thechapter.
 \ #1}{}}
\renewcommand{\chaptermark}[1]{
 \markboth{\uppercase{
 \chaptername\ \thechapter.
 \ #1}}{}}
\renewcommand{\chaptermark}[1]{
 \markboth{#1}{}}
\renewcommand{\chaptermark}[1]{
 \markboth{\thechapter.\ #1}{}}
\renewcommand{\chaptermark}[1]{
 \markboth{\thechapter.\
 \chaptername.\ #1}{}}

Prints:

Chapter 2. Do it now

CHAPTER 2. Do it now

CHAPTER 2. DO IT NOW

Do it now

2. Do it now

2. Chapter. Do it now

As an example let’s take a page layout where the leftmarks are generated by the sections and the rightmarks by the subsections (as is default in the article class). Take a page with some short sections, e.g. Section 1. subsection 1.1 subsection 1.2 Section 2.

As the leftmark contains the last mark of the page it will be “Section 2.”, and the rightmark will be “subsection 1.1” as it will be the first mark of the page. So the page header info will combine section 2 with subsection 1.1 which isn’t very nice. The best you can do in these cases is use only the \rightmarks and redefine \sectionmark accordingly. A L^{A}TEX command \firstleftmark would also be a nice addition (see the extramarks package in section 16).

Another problem with the marks in the standard L^{A}TEX classes is that the higher level sectioning commands (e.g. \chapter) call \markboth with an empty right argument. This means that on the first page of a chapter (or a section in article style) the \rightmark will be empty. If this is a problem you must manually insert extra \markright commands or redefine the \chaptermark (\sectionmark) commands to issue a \markboth command with two decent parameters.

As a final remark you should also note that the * forms of the \chapter etc. commands do not call the mark commands. So if you want your preface to set the header info but not be numbered nor be put in the table of contents, you must issue the \markboth command yourself, e.g.

\chapter*{Preface\markboth{Preface}{}}

Entering the \markboth command inside the \chapter* insures that the mark will not be separated from the title by a page break. Of course with \chapter this wouldn’t be a problem if you put the mark command after the chapter title, as the \chapter command starts at a new page. However with a \section* it could be dangerous to say:

\section*{Preface}
\markboth{Preface}{}

as a page break may occur between the two commands.

9 Dictionary style headers

Dictionaries and concordances usually have a header containing the first word defined on the page or both the first and the last words. This can easily be accomplished with fancyheadings and L^{A}TEX’s mark mechanism. Of course if you use the marks for dictionary style headers, you cannot use them for chapter and section information, so if there are also chapters and sections present, you must redefine the \chaptermark and \sectionmark to make them harmless:

\renewcommand{\chaptermark}[1]{}
\renewcommand{\sectionmark}[1]{}

Now you do a \markboth{#1}{#1} for each dictionary or concordance entry #1 and use \rightmark for the first entry defined on the page and \leftmark for the last one.

If you want to use a header entry of the form firstword–lastword it would be nice if this would be reduced to just the form firstword if both are the same. This could happen if there is just one entry on the page. In this case a test must be made to check if the marks are the same. However, TEX’s marks are strange beasts, which cannot be compared out of the box with the plain TEX \if commands. Fortunately the ifthen package works well:

\newcommand{\mymarks}{
  \ifthenelse{\equal{\leftmark}{\rightmark}}
    {\rightmark}
    {\rightmark--\leftmark}}
\lhead[\rm \mymarks]{\rm \thepage}
\rhead[\rm \thepage]{\rm \mymarks}

Dictionaries are often done with two columns. Unfortunately there is a bug in L^{A}TEX’s twocolumn option which causes some marks to be lost. If you use the package in figure 2, this will be solved^{2}^{3}.

2This is a simplification of the actual code. Get the real package from the CTAN sites.

3The multicol package uses a similar technique.

Figure 2 Fix marks in twocolumn style

% fixmarks.sty:
% Patch LaTeX's output routine to handle marks correctly with two columns.
% Joe Pallas <pallas@edu.stanford.neon>
% Corrected by Piet van Oostrum <piet@cs.ruu.nl> on Feb 5, 1993, Oct 5, 1994
\def\@outputdblcol{\if@firstcolumn \global\@firstcolumnfalse
% Remember the marks from the first column
    \global\setbox\@leftcolumn\copy\@outputbox
    \splitmaxdepth=\maxdimen \cbaddness=10000
    \setbox\@outputbox\vsplit\@outputbox to\maxdimen
    \xdef\@firstcoltopmark{\topmark}%
    \xdef\@firstcolfirstmark{\splitfirstmark}%
    \ifx\@firstcolfirstmark\empty\global\let\@setmarks\relax\else
      \gdef\@setmarks{\let\firstmark\@firstcolfirstmark
        \let\topmark\@firstcoltopmark}%
    \fi
% End of change
  \else \global\@firstcolumntrue
    \setbox\@outputbox\vbox{\hbox to\textwidth{\hbox to\columnwidth
      {\box\@leftcolumn \hss}\hfil \vrule width\columnseprule\hfil
       \hbox to\columnwidth{\box\@outputbox \hss}}}\@combinedblfloats
% Override current first and top with those of first column if necessary
       \@setmarks
% End of change
       \@outputpage \begingroup \@dblfloatplacement \@startdblcolumn
       \@whilesw\if@fcolmade \fi{\@outputpage\@startdblcolumn}\endgroup
    \fi}

10 Fancy layouts

You can make a multi-line field with the \\ command. It is also possible to put extra space in a field with the \vspace command. Note that if you do this you will probably have to increase the height of the header (\headheight) or of the footer (\footskip), otherwise you may get error messages “Overfull \vbox … has occurred while \output is active”. See Section 4.1 of the L^{A}TEX Companion for detail.

For instance, the following code will place the section title and the subsection title of an article in two lines in the upper right hand corner:

\documentclass{article}
\usepackage{fancyheadings}
\pagestyle{fancy}
\addtolength{\headheight}{\baselineskip}
\renewcommand{\sectionmark}[1]{\markboth{#1}{}}
\renewcommand{\subsectionmark}[1]{\markright{#1}}
\rhead{\leftmark\\\rightmark}

You can customize the decorative lines. You can make the decorative line in the header quite thick with

\setlength{\headrulewidth}{0.6pt}

or you can make the decorative line in the footer disappear with

\setlength{\footrulewidth}{0pt}

The decorative lines, themselves, are defined in the two macros \headrule and \footrule. For instance, if you want a dotted line rather than a solid line in the header, redefine the command \headrule:

\renewcommand{\headrule}{\vbox to 0pt{\hbox
    to\headwidth{\dotfill}\vss}}

11 Two book examples

The following definitions give an approximation of the style used in L. Lamport’s L^{A}TEX book.

Lamport’s header overhangs the outside margin. This is done as follows. The width of headers and footers is \headwidth, which by default equals the

width of the text: \textwidth. You can make the width wider (or narrower) by

redefining \headwidth with the \setlength and \addtolength commands. To

overhang the outside margin where the marginal notes are printed, add both

\marginparsep and \marginparwidth to \headwidth with the commands:
\addtolength{\headwidth}{\marginparsep}
\addtolength{\headwidth}{\marginparwidth}

You must issue these commands after the first \pagestyle{fancy} or \pagestyle{fancyplain} command as this will establish the default for

\headwidth.

And now a complete definition of Lamport’s book style:

\documentclass{book}
\usepackage{fancyheadings}
\pagestyle{fancyplain}
\addtolength{\headwidth}{\marginparsep}
\addtolength{\headwidth}{\marginparwidth}
\renewcommand{\chaptermark}[1]{\markboth{#1}{}}
\renewcommand{\sectionmark}[1]{\markright{\thesection\ #1}}
\lhead[\fancyplain{}{\bfseries\thepage}]
   {\fancyplain{}{\bfseries\rightmark}}
\rhead[\fancyplain{}{\bfseries\leftmark}]
   {\fancyplain{}{\bfseries\thepage}}
\cfoot{}

Notice that the \chaptermark and \sectionmark commands have been redefined to eliminate the chapter numbers and the uppercaseness.

For the second example, we take the AMS-L^{A}TEX book.

Chapter pages have no headers or footers. So we declare

\thispagestyle{empty}

for every chapter page, and we do not need fancyplain.

Chapter and section titles appear in the form: 2. DO IT NOW, so we have to redefine \chaptermark and \sectionmark as follows (see Section 8):

\renewcommand{\chaptermark}[1]%
   {\markboth{\uppercase{\thechapter.\ #1}}{}}
\renewcommand{\sectionmark}[1]%
   {\markright{\uppercase{\thesection.\ #1}}}

In an even-header, the page number is printed as the LeftHeader and the chapter info as the RightHeader; in an odd-header, the section info is printed as the LeftHeader and the page number as the RightHeader. The CenteredHeaders are empty. There are no footers.

There is a decorative line in the header. It is 0.5pt wide, so we need the commands:

\setlength{\headrulewidth}{0.5pt}
\setlength{\footrulewidth}{0pt}

The font used in the headers is 9 pt bold Helvetica. The PSNFSS system by Sebastian Rahtz uses the short (Karl Berry) name phv for Helvetica, so this font is selected with the commands:

\fontfamily{phv}\fontseries{b}\fontsize{9}{11}\selectfont

(See Sections 7.6.1 and 11.9.1 of the L^{A}TEX Companion.) Let us define a shorthand for this:

\newcommand{\helv}{%
   \fontfamily{phv}\fontseries{b}\fontsize{9}{11}\selectfont}

Now we are ready for the page layout:

\lhead[\helv \thepage]{\helv \rightmark}
\rhead[\helv \leftmark]{\helv \thepage}
\cfoot{}

Putting this all together:

\documentclass{book}
\usepackage{fancyheadings}
\pagestyle{fancy}
\renewcommand{\chaptermark}[1]%
   {\markboth{\uppercase{\thechapter.\ #1}}{}}
\renewcommand{\sectionmark}[1]%
   {\markright{\uppercase{\thesection.\ #1}}}
\setlength{\headrulewidth}{0.5pt}
\setlength{\footrulewidth}{0pt}
\newcommand{\helv}{%
   \fontfamily{phv}\fontseries{b}\fontsize{9}{11}\selectfont}
\lhead[\helv \thepage]{\helv \rightmark}
\rhead[\helv \leftmark]{\helv \thepage}
\cfoot{}

12 Special page layout for float pages

Some people want to have a special layout for float pages (pages only containing floats). As these pages are generated autonomically by L^{A}TEX, the user doesn’t have any control over them. There is no \thispagestyle for float pages and any change of the page style will at least also affect the page before the float page. With fancyheadings you can give the following definition in your preamble^{4}:

\makeatletter
\def\iffloatpage#1#2{\if@fcolmade #1\else #2\fi}
\makeatother

In each of the header- or footer fields you can now specify

\iffloatpage{value for float page}{value for other pages}

You can even use this to get rid of the decorative line on float pages only by defining:

\renewcommand{\headrulewidth}{\iffloatpage{0pt}{0.4pt}}

Note that \renewcommand is used rather than the usual \setlength command to change the \headrulewidth. Changing it with \setlength will not work^{5}.

After this \renewcommand it will no longer be possible to change the param-

eter \headrulewidth with \setlength, you will have to use \renewcommand in

the rest of the document.

Sometimes you may want to change the layout also for pages that contain a float on the top of the page or a float on the bottom of the page.

Enter the following definitions in a style file (or between \makeatletter and \makeatother in your preamble:

\let\latex@makecol\@makecol
\def\@makecol{\let\topfloat\@toplist\let\botfloat\@botlist\latex@makecol}
\def\iftopfloat#1#2{\ifx\topfloat\empty #2\else #1\fi}
\def\ifbotfloat#1#2{\ifx\botfloat\empty #2\else #1\fi}

You can then use the commands \iftopfloat and \ifbotfloat similar to

\iffloatpage.

Note: Marks in floats will not be visible in L^{A}TEX’s output routine, so it is not useful to put marks in floats. So there is currently no way to let a float (e.g. a figure caption) influence the page header or footer.

13 Those blank pages

In the book class when the openany option is not given or in the report class when the openright option is given, chapters start at odd-numbered pages, half of the time causing a blank page to be inserted. Some people prefer this page to be completely empty, i.e. without headers and footers. This cannot be done with \thispagestyle as this command would have to be issued on the previous page. There is, however, no magic necessary to get this done:

4If you have fancyheadings version 1.5 or higher the commands described in this section are already included

5It was probably a dumb decision to define the width parameters as lengths. There is no need to and I may change this in a future version of fancyheadings

\clearpage{\pagestyle{empty}\cleardoublepage}

As the \pagestyle{empty} is enclosed in a group it only affects the page that may be generated by the \cleardoublepage. You can of course put the above in a private command. If you want to have this done automatically at each chapter start or when you want some other text on the page then you must redefine the \cleardoublepage command.

\makeatletter
\def\cleardoublepage{\clearpage\if@twoside \ifodd\c@page\else
  \hbox{}
  \vspace*{\fill}
  \begin{center}
    This page intentionally contains only this sentence.
  \end{center}
  \vspace{\fill}
  \thispagestyle{empty}
  \newpage
  \if@twocolumn\hbox{}\newpage\fi\fi\fi}
\makeatother

14 N of M style page numbers

Some document writers prefer the pages to be numbered as n of m where m is the number of pages in the document. There is a package nofm.sty available, but some versions of it are defective, and most don’t work with fancyheadings because they take over the complete page layout. For L^{A}TEX 2ε there is a package lastpage available which you can use with fancyheadings as follows:

\usepackage{lastpage}
...
\cfoot{\thepage\ of \pageref{LastPage}}

If you are still using L^{A}TEX2.09 and you are not able to switch to L^{A}TEX 2ε you can define a L^{A}TEX2.09 compatible lastpage.sty as follows:

\let\origenddocument=\enddocument
\def\enddocument{\clearpage\if@filesw
  {\addtocounter{page}{-1} \immediate\write\@mainaux
    {\string\newlabel{LastPage}{{}{\thepage}}}}\origenddocument}

With some trickery^{6} the value of the LastPage label can be used to make different headers or footers on the last page of a document. E.g. if you want the footer of every odd page, except if it is the last one, to contain the text “please turn over”, this can be done as follows:

6If you have a recent version of the ifthen package the tricks are not necessary. You can then test the \pageref{LastPage} directly.

\usepackage{lastpage}
\usepackage{ifthen}
...
\makeatletter
\def\LastPagenum{\@ifundefined{r@LastPage}{0}{\expandafter
    \expandafter\expandafter\@cdr
    \csname r@LastPage\endcsname\@nil\null} }
\makeatother
\rfoot{\ifthenelse{\isodd{\value{page}} \and \not
    \value{page}=\LastPagenum}{please turn over}{}}

15 When to change the headers and footers?

Sometimes you want to change the header or footer layout in the course of a document. Some of these changes can be accomplished by using the mark mechanism as may be seen in section 8 and 16. However, sometimes we want a more drastic change, e.g to change the page numbering from roman to arabic (with \pagenumbering), to change one of the fancyheadings fields or to change to another page style. Sometimes you may be surprised to find the change to occur too early. In general the above mentioned changes take effect immediately, i.e. on the page that is currently being built. If you want the change to take effect at the next page you must make sure that the current page is finished. In most cases this can be done by issuing a \clearpage command before any of the above mentioned changes. If this is not possible you can use the afterpage package with:

\afterpage{\lhead{new value}} or \afterpage{\pagenumbering{roman}}.

You cannot use \afterpage to change the \pagestyle as the commands issued by \afterpage are local in a group, and the \pagestyle command makes only local changes. The \pagenumbering and the fancyheadings commands make global changes so they will work, as will the \thispagestyle command.

It should be noted that although the fancyheadings commands like \lhead take effect immediately, this does not mean that any “variables” used in these commands get the value they have at the place where these commands are given. E.g. if \cfoot{\thepage} is given the page number that will be inserted in the footer is not the page number of the page where this command is given, but rather the page number of the actual page where the footer is constructed. Of course for the page number this is what you expect, but it is also true for other commands.

So if you have a book where each chapter is written by a different author and you want the name of the author in the lower left-hand corner you can use the following commands:

\newcommand{\TheAuthor}{}
\newcommand{\Author}[1]{\renewcommand{\TheAuthor}{#1}}
\lfoot{\TheAuthor}

and start each chapter with the command \Author{Real Name}. If however, the author name would be changed before a page is completed the wrong author could come in the footer. This would be the case if you gave the above command before the \chapter command rather than after it. Another source of problems is the fact that TEX’s output routine processes commands ahead, so it may already have processed some commands that produce text that will appear on the next page. See the next section for an example.

16 Headers and footers induced by the text

We have seen how we can use L^{A}TEX’s marks to get information from the document contents to the headers and footers. The marks mechanism is the only reliable mechanism that you can use to get changing information to the headers or footers. This is because L^{A}TEX may be processing your document ahead before deciding to break the page.

Sometimes the two marks that L^{A}TEX offers are not enough. An example is the following:

If a solution to an exercise goes across a page break, then I would like to have “(Continued on next page…)” at the bottom of the first page and “(Continued…)” at the top in the margin of the next page.

You cannot use L^{A}TEX’s mark mechanisms for this if you also want to use chapter and section information.

The code from figure 3 constitutes a package that gives you two extra marks that can be used in this situation^{7}. Here is a way to use this package:

\usepackage{extramarks}
...
\pagestyle{fancy}
\lhead{\firstxmark}
\rfoot{\lastxmark}
...
\extramarks{}{Continued on next page\ldots}
Some text that may or may not cross a page boundary...
\extramarks{Continued\ldots}{}

Note that the \extramarks command must be close to the text, i.e no empty lines (paragraph boundaries) should intervene. Otherwise the page may be broken at that boundary and the extramarks would come on the wrong page.

There are two new marks that can be used in the page layout with this package: If commands of the form \extramarks{m_{1}}{m_{2}} are given \firstxmark gives you the first m_{1} value and \lastxmark gives you the last m_{2} value of the current page. It also gives you the \firstleftmark and \lastrightmark commands that complement the standard L^{A}TEX marks.

To stress the point that marks are the correct way to do this, let me give you a “solution” that will not work^{8}:

7After I made this package I discovered a package secret.sty that does a similar thing to mark confidential paragraphs if they cross a page boundary. It does it, however, by changing the output routine.

8Actually there is another way but it requires two L^{A}TEX passes: you can put \label commands before and after the text and compare the \pagerefs.

Figure 3 Package for extra marks in L^{A}TEX

% extramarks.sty
\def\@leftmark#1#2#3#4{#1}
\def\@rightmark#1#2#3#4{#2}
\def\markboth#1#2{{\def\protect{\noexpand\protect\noexpand}
     \let\label\relax \let\index\relax \let\glossary\relax
     \expandafter\@markboth\@themark{#1}{#2}
     \mark{\@themark}}\if@nobreak\ifvmode\nobreak\fi\fi}
\def\markright#1{{\def\protect{\noexpand\protect\noexpand}
     \let\label\relax \let\index\relax \let\glossary\relax
     \expandafter\@markright\@themark
     {#1}\mark{\@themark}}\if@nobreak\ifvmode\nobreak\fi\fi}
\def\@markright#1#2#3#4#5{\gdef\@themark{{#1}{#5}{#3}{#4}}}
\def\@markboth#1#2#3#4#5#6{\gdef\@themark{{#5}{#6}{#3}{#4}}}
\def\leftmark{\expandafter\@leftmark\botmark{}{}{}{}}
\def\rightmark{\expandafter\@rightmark\firstmark{}{}{}{}}
\def\firstleftmark{\expandafter\@leftmark\firstmark{}{}{}{}}
\def\lastrightmark{\expandafter\@rightmark\botmark{}{}{}{}}
\def\@themark{{}{}{}{}}
\def\extramarks#1#2{{\def\protect{\noexpand\protect\noexpand}
     \let\label\relax \let\index\relax \let\glossary\relax
     \expandafter\@markextra\@themark{#1}{#2}
     \mark{\@themark}}\if@nobreak\ifvmode\nobreak\fi\fi}
\def\@markextra#1#2#3#4#5#6{\gdef\@themark{{#1}{#2}{#5}{#6}}}
\def\firstxmark{\expandafter\@firstxmark\firstmark{}{}{}{}}
\def\topxmark{\expandafter\@firstxmark\topmark{}{}{}{}}
\def\lastxmark{\expandafter\@lastxmark\botmark{}{}{}{}}
\def\@firstxmark#1#2#3#4{#3}
\def\@lastxmark#1#2#3#4{#4}

\lhead{Continued}
\rfoot{Continued on next page\ldots}
Some text that may or may not cross a page boundary...
\lhead{}
\rfoot{}

You may be tempted to think that the first \lhead and \rfoot will be in effect when TEX breaks the page in the middle of the text, and the last ones when the page breaks after the text. This is not true as the whole paragraph (including the last definitions) will be processed before TEX considers the page break, so at the time of the page break the last definitions are effective, whether the page break occurs inside the text or outside of it. Putting a paragraph boundary between the text and the last definitions will not work either, because you don’t want the first definitions to be in effect when TEX decides to break the page exactly at this boundary. Actually the marks mechanism was invented to get rid of these kinds of problems.

In the above example the text “Continued” appears in the page header. It may be nicer to put it in the margin. This can be easily accomplished by positioning it at a fixed place relative to the page header. In plain TEX you

would use a concoction of \hbox to 0pt, \vbox to 0pt, \hskip,\vskip, \hss

and \vss but fortunately L^{A}TEX’s picture environment gives a much cleaner way to do this. In order not to disturb the normal header layout we put the text in a zero-sized picture. Generally this is the best way to position things on fixed places on the page. You can then also use the normal headings. See also section 18 for another example of this technique.

\setlength{\unitlength}{\baselineskip}
\lhead{\begin{picture}(0,0)
  \put(-2,-3){\makebox(0,0)[r]{\firstxmark}}
  \end{picture}\leftmark}

This solution can of course also be used for the footer. Be sure to put the picture as the first thing in left-handside entries and last in right-handside ones.

Finally you may want to put “(Continued…)” in the text rather than in the header or the margin. Then you have to use the afterpage package. We also decide to make a separate environment for it.

\newenvironment{continued}{\par
  \extramarks{}{Continued on next page\ldots}
  \afterpage{\noindent\firstxmark\vspace{1ex}}
  }{\extramarks{(Continued\ldots)}{}\par}

It is a bit dangerous to use \firstxmark outside the page layout routine, but apparently with \afterpage this works. If you would need the information further on in the page you must remember the state of the marks in your own variable. You can do this in one of the fancyheadings fields. For example if you want to add something after the broken piece of text you can use the following:

\newcommand{\mysaved}{}
\newenvironment{continued}{\par
  \extramarks{}{Continued on next page\ldots}
  }{\extramarks{(Continued\ldots)}{}\par\vspace{1ex}\mysaved}
\lhead{\leftmark}
\chead{\ifthenelse{\equal{\lastxmark}{}}
  {\gdef\mysaved{}}
  {\gdef\mysaved{\noindent[Continued from previous page]}}}

If you want to include one of the marks or other varying information in the saved text, you must use \xdef rather than \gdef.

17 A movie

If you put at each page on the same place a picture that slightly changes from page to page you can get a movie-like effect by flipping through the pages. You can create such a movie easily with fancyheadings. For simplicity we assume that each picture is in a postscript (EPS) file called pic⟨n⟩.ps where ⟨n⟩ is the page number and that we use the epsf package. To put the movie in the righthandside bottom corner the following will work:

\rfoot{\setlength{\unitlength}{1mm}\setlength{\epsfxsize}{2cm}
  \begin{picture}(0,0)
    \put(5,0){\epsfbox{pic\thepage.ps}}
  \end{picture}}

Notice that the \unitlength and \epsfxsize parameter should be set locally in the fancyheadings field in order to avoid unwanted interference from their values in the text.

18 Thumb-indexes

Some railroad guides and expensive bibles have so called thumb-indexes, i.e. there are marks on the sides of the pages that indicate where the chapters are. You can create these by printing black blobs in the margin of the pages. The vertical position should be determined by the chapter number or some other counter. As the position is independent of the contents of the page, we print these blobs as part of the header in a zero-sized picture as described in the previous section.

Of course we have to take care of two-sided printing, and we may want to have an index page with all the blobs in the correct position. The solution requires some hand-tuning to get the blobs nicely spaced out vertically. For the application that I had there were 12 sections, so I made the blobs 18 mm apart, i.e. 9 mm blob separated by 9 mm whitespace. In order to avoid calculations they are set in a picture environment with the \unitlength set to 18 mm. Page numbers are set in the headers at the outer sides, and the blobs are attached to these. In this example the section numbers are used to position the blobs, but you can replace this with any numeric value. See figure 4 for the resulting overview page and figure 5 for the code.

Figure 4 Thumb-index overview page

1_{1 Introduction 2 The first year 3 Specialisation}

19 Final Remarks

The commands to specify the headers and footers are quite complicated when even- and odd-numbered pages have different layout and we want also to redefine the plain page style. Therefore for version 2 of fancyheadings I will create Figure 5 Thumb-index code

\setlength{\unitlength}{18mm}
\newcommand{\blob}{\rule[-.2\unitlength]{2\unitlength}{.5\unitlength}}
\newcommand\rblob{\thepage
  \begin{picture}(0,0)
    \put(1,-\value{section}){\blob}
  \end{picture}}
\newcommand\lblob{%
  \begin{picture}(0,0)
    \put(-3,-\value{section}){\blob}
  \end{picture}%
  \thepage}
\pagestyle{fancy}
\cfoot{}
\newcounter{line}
\newcommand{\secname}[1]{\addtocounter{line}{1}%
  \put(1,-\value{line}){\blob}
  \put(-7.5,-\value{line}){\Large \arabic{line}}
  \put(-7,-\value{line}){\Large #1}}
\newcommand{\overview}{\thepage
  \begin{picture}(0,0)
    \secname{Introduction}
    \secname{The first year}
    \secname{Specialisation}
...etc...
  \end{picture}}
\begin{document}
\rhead{\overview}\mbox{}\newpage % This produces the overview page
\rhead{} % Front matter may follow here
\clearpage
\rhead[\rightmark]{\rblob} \lhead[\lblob]{\leftmark}
...

an easier user interface (the old commands will still work). To give you the opportunity to comment, here is a preview:

The *head and *foot commands will be replaced by commands \fancyhead and \fancyfoot. These will have an optional argument that selects which fields to change, e.g. L for the left field, E for even pages, etc. See figure 6. Selectors can be combined so \fancyhead[LE,RO]{text} will define the field for both the left header on even pages and the right header on odd pages. There will also be a new command \fancypagestyle that allows you to (re)define page styles. One use for this is to redefine the page style plain so the \fancyplain stuff is no longer necessary. Another use is to allow special page styles to be used with thispagestyle or maybe for float pages.

The complicated example in section 6 will then become:

Figure 6 Selectors

E Even page
O Odd page
L Left field
C Center field
R Right field
T float at Top
B float at Bottom
F Float page

\fancyhead{} % clear all fields
\fancyhead[LE,RO]{\bfseries The performance of new graduates}}
\fancyfoot[LO,CE]{From: K. Grant}
\fancyfoot[RE,CO]{To: Dean A. Smith}
\fancyfoot[RO,LE]{\bfseries \thepage}
\setlength{\headrulewidth}{0.4pt}
\setlength{\footrulewidth}{0.4pt}
\fancypagestyle{plain}{%
\fancyhead{}
\fancyfoot[C]{\bfseries \thepage}}

As you can see this is much less complicated.

If you want to comment please use e-mail or regular mail to: Piet van Oostrum Dept. of Computer Science Utrecht University P.O. Box 80.089 3508 TB Utrecht, The Netherlands Telephone: +31 30 531806 Telefax: +31 30 513791 Internet: piet@cs.ruu.nl

Index

afterpage.sty, 15, 18

bible, 19 BIBLIOGRAPHY, 7 blank page, 14

\cfoot, 3
\chaptermark, 7, 8
\chaptername, 7
\chead, 3
\cleardoublepage, 14
\clearpage, 14, 15

concordance, 9 Continued…, 16

dictionary, 9

epsf, 18
\extramarks, 16, 17
extramarks.sty, 2
\firstleftmark, 8, 16
fixmarks.sty, 2, 10

float page, 13

\footrulewidth, 3, 6, 11
\footskip, 10

ftp, 2

\headheight, 10
\headrulewidth, 3, 6, 10
\headwidth, 11
\ifbotfloat, 13
\iffloatpage, 13
ifthen, 9
\iftopfloat, 13

INDEX, 7

lastpage.sty, 14
\lastrightmark, 16
\leftmark, 6, 7
\lfoot, 3
\lhead, 3
\maketitle, 3

margin, 17

\marginparsep, 11
\marginparwidth, 11

movie, 18 multi-line, 10

                                    nofm.sty, 14

page break, 17 page style

changes, 15

                                        empty, 14
                                        fancy, 2
                                        fancyplain, 4, 5, 20
                                        plain, 4, 20
                                    \pagenumbering, 15
                                    picture, 18
                                    \plainfootrulewidth, 5, 6
                                    \plainheadrulewidth, 5, 6
                                    \rfoot, 3
                                    \rhead, 3
                                    \rightmark, 6, 7
                                    \sectionmark, 7
                                    \subsectionmark, 7
                                    \thechapter, 7
                                    \thispagestyle, 3

thumb-index, 19

                                    twoside, 4
                                    \uppercase, 7, 8